OpenStudy (anonymous):

solve the given systems of equations using any method and then choose the correct ordered pair from the given choices: 2x + 3y = 6 5x + 6y = 9

OpenStudy (texaschic101):

Elimination : 2x + 3y = 6 --->(-2) 5x + 6y = 9 ------------ -4x - 6y = - 12 (result of multiplying by -2) 5x + 6y = 9 ------------add x = - 3 now sub -3 in for x in either equation to find y 2x + 3y = 6 2(-3) + 3y = 6 -6 + 3y = 6 3y = 6 + 6 3y = 12 y = 4 check.. 5x + 6y = 9 5(-3) + 6(4) = 9 -15 + 24 = 9 9 = 9 (correct) ordered pair that is a solution is : (-3,4)

OpenStudy (anonymous):

Thank you so much ! you tha real mvp haha . u wanna help me with another problem?

OpenStudy (texaschic101):

what ya got ?

OpenStudy (anonymous):

hold on